1-convert AB to base 10.
2-110101(base2)to base 10
3- Using an ASCII table, convert and answer the question below. Make sure to
answer in ASCII. 01010111 01101000 01100001 01110100 01101001
01110011 01110100 01110111 01101111 00101011 00110100 00111111
4- While working on a computer the following phrase pops up. What is going on
(hint use ASCII table)? 01010110 01101001 01110010
01110101 01110011
Answer #1

1) Given HexaDecimal AB

AB = (10 × 16¹) + (11 × 16⁰) = (171)10

2) Given Binary number is  110101

110101 = (1 × 2⁵) + (1 × 2⁴) + (0 × 2³) + (1 × 2²) + (0 × 2¹) + (1 × 2⁰) = (53)10

3) Given Binary number is 10101001

10101001 = (1 × 2⁷) + (0 × 2⁶) + (1 × 2⁵) + (0 × 2⁴) + (1 × 2³) + (0 × 2²) + (0 × 2¹) + (1 × 2⁰) = (169)10

Given Binmary Value is 010101110110100001100001011101000110100101110011011101000111011101101111001010110011010000111111
The ASCII Value is: Whatistwo+4?

Answer to the above ASCII Question is 6

4) Given Binmary Value is 0101011001101001011100100111010101110011
The Ascii Value is: Virus